### v2

# Electronic order induced symmetry breaking in lattice dynamics of Co3Sn2S2

Abstract
Based on the molecular Berry curvature (MBC) framework, we develop an ab initio algorithm to capture the quantitative effects of magnetic order on lattice dynamics. Using the ferromagnetic Weyl semimetal Co3Sn2S2 as a prototype, we show that electronic-order-driven phonon symmetry breaking requires spin-orbit coupling (SOC) and leads to an MBC term that breaks both time-reversal ($\mathcal{T}$) and mirror symmetries. We demonstrate that mirror-symmetry breaking is essential to account for the experimentally observed phonon splitting, $\mathcal{T}$-breaking alone is insufficient. The MBC is widely distributed across the Brillouin zone, giving rise to significant off-$\Gamma$ effects. Our results show the phonon splitting of the $E_{g}$ and $E_{u}$ modes have distinct origins: the $E_{g}$ splitting arises from MBC and is captured by our algorithm, while the $E_{u}$ splitting stems from correlation-sensitive Fano resonance, explaining their asymmetric experimental line shapes. We provide a design strategy and computational pathway for materials with large phonon magnetism and splitting, requiring SOC, electron-phonon coupling, and ideally flat bands at the Fermi level. This work also suggests new avenues for controlling non-reciprocal phonon transport.
Within Landau’s theory, electronic orders, such as magnetic/spin, charge, and orbital order, are typically characterized by local order parameters and strong electron interactions, and are invariably accompanied by spontaneous symmetry breaking. This contrasts with phonon systems, where such symmetry breaking is rare. Recent work has shown that time-reversal symmetry ($\mathcal{T}$) breaking occurs in phonons by coupling to external fields [1, 2, 3, 4], spins [5, 6] and chiral excitations [7]. Moreover, phonon energy splitting in ferromagnetic (FM) Weyl semimetal css directly demonstrates this effect [8, 9]. Although ab initio methods can readily describe electronic symmetry breaking, such as $\mathcal{T}$ in magnetic systems, the concomitant symmetry breaking in the phonon system induced by electronic order remains inadequately explored by first-principles calculations. Thus, a quantitative understanding of symmetry breaking in lattice dynamics from first principles [10, 11, 12] is essential, and it is crucial for interpreting phonon-related transport [13, 14, 15, 16, 17, 15, 18, 19, 20, 4] and optical properties [9, 8, 21, 22, 1].
To capture the experimentally observed phonon splitting in css, which is induced by magnetic order, we develop an ab initio algorithm based on molecular Berry curvature (MBC, or nuclear Berry curvature) theory. A nonzero MBC contribution necessitates the inclusion of spin-orbit coupling (SOC) [10]. This requirement distinguishes our approach from the phonon self-energy method in Ref. [11], which yields phonon splitting even in the absence of SOC. Under FM order and SOC, the MBC contribution breaks both $\mathcal{T}$ and mirror symmetries in css. Our results demonstrate that this mirror-symmetry breaking is essential to explain the observed phonon splitting in css, as $\mathcal{T}$ breaking alone is insufficient [23]. Furthermore, there’s a broad Brillouin zone (BZ) distribution of the MBC contributions, leading to significant effects away from the $\Gamma$-point. Both mirror-symmetry breaking and off-$\Gamma$ effects, which remain largely unexplored, are critical for a complete theoretical description of the lattice dynamics in css. Furthermore, our results demonstrate that the magnitude of the phonon splitting, while relevant for circularly polarized phonons with nonzero angular momentum (i.e., chiral phonons), is not uniquely determined by its absolute value.
Figure 1: The crystal structure, Brillouin zone, electronic/phonon structure of css. (a) Crystal structure of css. Brown arrows indicate the local magnetic moments on Co atoms. This magnetic order breaks the vertical mirror and time-reversal symmetry, but keeps the $C_{3}$ and inversion symmetry. (b) Brillouin zone of css. (c) Electronic band structure with FM order and spin-orbit coupling. (d) Phonon spectrum of css without the molecular Berry curvature contribution. The coordinates of $K$ are ($\frac{1}{3}$, 0, -$\frac{1}{3}$)/ ($\frac{1}{3}$, $\frac{1}{3}$, 0) in the basis of the primitive/conventional lattice vectors.
Lattice dynamics for css without MBC css is the first experimentally confirmed magnetic Weyl semimetal [24, 25, 26, 27, 28, 29, 30, 31, 32, 33] with a centrosymmetric space group of $R\bar{3}m$ ($\mathcal{P}$-preserving), thus the phonon modes can be classified in Raman active ones and infrared (IR) active ones at $\Gamma$. Figure 1(a) shows the primitive cell of css with brown arrows indicating the local magnetic moment. Below $T_{\rm C}\sim 175$[34, 27, 28, 29, 35, 36, 37, 38], FM order with $C_{3}$ axis spin alignment breaks both $\mathcal{T}$ and vertical mirror symmetry, enabling an SOC-essential Weyl semimetal phase, reducing the point group from $D_{3d}$ to $C_{3i}$ (magnetic space group $R\bar{3}m^{\prime}$).
Figure 1(c) presents the electronic band structure of css with FM order and SOC. The Fermi level has been shifted by 25 meV to align with the in-plane flat band position observed by the scanning tunneling microscopy [39]. The electronic bands near the Fermi level are primarily contributed by the $d$-orbitals of Co atoms.
Figure 1(d) displays the phonon spectrum without the MBC contribution. The highlighted $E_{u}$ and $E_{g}$ modes correspond to those observed experimentally [9, 8], and will be discussed in detail through quantitative comparisons. The sole $E_{g}$ mode arises from S atoms occupying the $6c$ Wyckoff positions, which lacks inversion site-symmetry. In contrast, Co, Sn1 and Sn2 reside at the Wyckoff positions of $9e$, $3a$ and $3b$, respectively, all of them preserve inversion site-symmetry. The atomic displacements transform as a vector representation under $O(3)$ operations, analogous to $p$-orbitals, and exhibit odd parity. As a result, neither Co nor Sn can generate phonon modes of even parity.
The degeneracy of the $E_{g}$ and $E_{u}$ modes along the $C_{3}$-invariant path, i.e., $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$, is protected by vertical mirror symmetry as well as $\mathcal{PT}$ symmetry [23]. The experimental observation of phonon splitting thus implies the breaking of both $\mathcal{PT}$ and vertical mirror symmetry, as shown in Fig. 2(e). To quantitatively describe the underlying physical mechanism, we develop a ab initio+MBC algorithm based on the Wannier basis to efficiently quantify this ferromagnetism-induced $E_{g/u}$ splitting.
Figure 2: Phonon spectra near the $\Gamma$ point under successive symmetry-breaking conditions. From (a) to (c), vertical mirror symmetry ($\sigma_{v}$) and inversion symmetry ($\mathcal{P}$) are progressively broken. From top to bottom, time-reversal symmetry ($\mathcal{T}$) is broken. Since css belongs to the $D_{3d}$ point group, breaking $\mathcal{T}$ alone is insufficient to lift the phonon degeneracy, as seen between (a) and (d). Therefore, the MBC mechanism, which involves spin–orbit coupling and breaks $\sigma_{v}$, is necessary to split the degenerate modes.
MBC modified lattice dynamics Experiments rule out resonant magnon-phonon coupling in css [40], allowing it to be safely discounted for an accurate description of the lattice dynamics [10, 2]. Furthermore, Ref. [9] confirms weak electron-hole pair excitation in Co3Sn2S2, preserving Born-Oppenheimer approximation (BOA) validity for MBC calculations. Within the BOA, the adiabatic evolution of the electronic ground state with the lattice accumulates a Berry phase, which subsequently couples back to influence lattice dynamics [41, 20], thus the equation of motion (EOM) reads:
$\tilde{K}(\bm{q})+\tilde{G^{\dagger}}(\bm{q})\tilde{G}(\bm{q})]\epsilon_{\nu\bm{q}}=\omega^{2}_{\nu\bm{q}}\epsilon_{\bm{q}}-2i\omega_{\bm{q}}\tilde{{G}}(\bm{q})\epsilon_{\nu\bm{q}}$}, (1)
where $\epsilon_{\nu\bm{q}}$ is the polarization vector, $\tilde{K}_{\kappa\alpha,\kappa^{\prime}\beta}(\bm{q})=\frac{1}{\sqrt{M_{\kappa}M_{\kappa^{\prime}}}}K_{\kappa\alpha,\kappa^{\prime}\beta}(\bm{q})$ and $\tilde{G}_{\kappa\alpha,\kappa^{\prime}\beta}(\bm{q})=\frac{1}{2\sqrt{M_{\kappa}M_{\kappa^{\prime}}}}G_{\kappa\alpha,\kappa^{\prime}\beta}(\bm{q})$ are the mass-weighted dynamic matrix and MBC. To solve this equation, we can take $\tilde{G}(\bm{q})$ as the first-order perturbation, and discard the second-order perturbation $\tilde{G^{\dagger}}(\bm{q})\tilde{G}(\bm{q})$. The zero-order equation reads:
$\tilde{K}(\bm{q})\epsilon^{(0)}_{\nu\bm{q}}=\omega^{(0)2}_{\nu\bm{q}}\epsilon^{(0)}_{\bm{q}}$}, (2)
and the first-order modification of $\omega^{(0)2}_{\nu\bm{q}}$ reads
$\Delta\omega^{(0)2}_{\nu\bm{q}}=\langle\epsilon_{\nu\bm{q}}|2i\omega^{(0)}_{\nu\bm{q}}\tilde{G}_{\kappa\alpha,\kappa^{\prime}\beta}(\bm{q})|\epsilon_{\nu^{\prime}\bm{q}}\rangle.$ (3)
$\nu=\nu^{\prime}$ for the non-degenerated case. Detailed discussions are in the Supplementary Material.
Standard first-principles calculations of the dynamical matrix $K_{ij}$ preserve the symmetries that are broken by the electronic FM order parameterized via a pseudovector, such as $\mathcal{T}$ and vertical mirror symmetry (see Supplementary Material). Thus, this algorithmic limitation results in a spurious double degeneracy of the $E_{g}$ and $E_{u}$ phonon modes, as shown in Fig. 2(a). Although the MBC term $\tilde{G}$ was proposed to break $\mathcal{T}$-symmetry [20, 20, 10], it alone is insufficient to lift the degeneracy of phonon bands along the $C_{3v}$-invariant high-symmetry path ($\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$) in css, due to the presence of vertical mirror symmetry [23], as illustrated in Fig. 2 (d). Therefore, an additional mechanism breaking the both of the $\mathcal{T}$ and vertical mirror symmetry by MBC is essential to explain the experimentally observed phonon splitting in Fig. 2 (e).
Figure 3: The MBC induced splitting for $E_{g}$ mode. (a) Phonon dispersion of the $E_{g}$ mode along the $C_{3}$-invariant path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$, calculated without (black dashed line) and with molecular Berry curvature (blue and red solid lines). The colored stars mark the energy positions of the $E_{g}$ mode observed experimentally. (b) Energy splitting of the $E_{g}$ mode along the $C_{3}$-invariant path. The insert figure shows the vibrational modes of the S atoms at the $\Gamma$ and T points. (c) Evolution of the pseudo-angular momentum ($l_{ph}$) and the $z$-component of the angular momentum ($l_{z,\nu\bm{q}}$) along the $C_{3}$-invariant path. (d) Atomic displacement patterns of the two $E_{g}$ modes at the $\Gamma$ point.
MBC-modified $E_{g}$ mode Figure 3(a) displays the phonon spectra of the $E_{g}$ mode with MBC along the $C_{3}$-invariant path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$. Within our ab initio framework, MBC lifts the degeneracy along the entire path by breaking both time-reversal $\mathcal{T}$ (as well as $\mathcal{PT}$) and vertical mirror symmetries. The magnitude of the phonon energy splitting for the $E_{g}$ modes is presented in Fig. 3(b), revealing a maximum splitting at the $\Gamma$ point and a minimum splitting at T.
At $\Gamma$, the calculated phonon splitting is approximately 0.253 meV (2.05 cm-1), which agrees with the experimental value of 1.27 cm-1 reported in Ref. [8] in order of magnitude. The minor discrepancy may originate from slight variations in the experimental Fermi level and anharmonic effects [42, 43]. To further investigate the influence of the Fermi level on the MBC, we shifted it by 60 meV and recalculated the splitting of the $E_{g}$ mode. This adjustment yields a splitting of approximately 1.24 cm-1, which is in excellent agreement with the experimental value (see Supplementary Material for details).
To further elucidate the lattice dynamics of $E_{g}$ induced by MBC, we calculate both the pseudo-angular momentum ($l_{ph}$) and the $z$-component of the angular momentum ($l_{z,\nu\bm{q}}$) along the $C_{3}$-invariant path, as shown in Fig. 3(c). We note the angular momentum only has $z$ component due to the restriction of $C_{3}$ symmetry [23]. Both quantities exhibit a quantized value of unity, with sign reversals resulting from band inversion along the path, which fundamentally governs the emergence of the phonon Hall effect [13, 14, 15, 16, 17, 15, 18, 19, 20, 4].
The quantized unity of the $l_{\text{ph}}$ arises from the $C_{3}$ symmetry of the system. In contrast, the quantization of $l_{z,\nu\bm{q}}$ along the path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$ can be attributed to two factors: (i) At $\Gamma$, $E_{g}$ is contributed solely by the vibration of S atoms residing at $C_{3}$-invariant Wyckoff positions [44, 23]. As illustrated in Fig. 3(d), these two S atoms rotate in the same direction but with a $\pi$ phase difference at the $\Gamma$. (ii) Away from $\Gamma$, their energies are well-separated from other modes, resulting in minimal hybridization with other phonon modes.
It should be noted that the phonon splitting induced by MBC is not uniquely determined by $l_{z,\nu\bm{q}}$. As shown in Figs. 3(b) and (c), $l_{z,\nu\bm{q}}$ remains close to unity along the $C_{3}$-invariant path, yet the splitting of the $E_{g}$ modes varies significantly with a maximum at $\Gamma$.
Figure 4: The MBC induced splitting for $E_{u}$ mode. (a) Phonon dispersion of the $E_{u}$ mode along the $C_{3}$-invariant path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$, calculated without (black dashed line) and with molecular Berry curvature (blue and red solid lines). The colored stars mark the energy positions of the $E_{u}$ mode observed experimentally. (b) Energy splitting of the $E_{u}$ mode along the $C_{3}$-invariant path. The insert figure shows the vibrational modes of the S atoms at the $\Gamma$ and T points. (c) Evolution of the pseudo-angular momentum ($l_{\text{ph}}$) and the $z$-component of the angular momentum ($l_{z,\nu\bm{q}}$) along the $C_{3}$-invariant path. (d) Atomic displacement pattern of the two $E_{u}$ modes at the $\Gamma$ point.
MBC-modified $E_{u}$ mode Figure 4(a) displays the MBC-modified phonon spectra of the $E_{u}$ mode along the path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$. The MBC also lifts the degeneracy of the $E_{u}$ mode (dashed line) across the entire $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$ path. The magnitude of the phonon energy splitting for the Eg modes is presented in Fig. 4(b), revealing a maximum splitting in the vicinity of T and minimum splitting in the vicinity of $\Gamma$.
Combined with Figs. 3(b) and  4(b), it can be observed that although the parity of the $E_{g}$ and $E_{u}$ modes remains unchanged at both $\Gamma$ and T, the relative phase between the two S atoms shifts by $\pi$. For instance, the $E_{g}$ mode exhibits a relative phase of $\pi$ at $\Gamma$ but 0 at T, whereas the $E_{u}$ mode shows a phase of 0 at $\Gamma$ and $\pi$ at T. This switching of the relative phase between the two modes correlates with the reversal of the maximum phonon splitting.
At the $\Gamma$ point, the splitting calculated from our algorithm is about 0.002 meV (0.016 cm-1), which is smaller than the experimentally observed value of 0.3 cm-1 reported in the optical conductivity measurements [9]. This discrepancy suggests that the splitting observed in experiments is not primarily caused by the MBC. Combined with the asymmetric line shapes of the $E_{u}$ mode, we attribute the larger experimental splitting to a Fano resonance [9]. This resonance may arise from quantum interference between electronic and phonon excitations when Weyl cones approach the Fermi level [45, 46]. The absence of an asymmetric line shape in the $E_{g}$ mode further supports this interpretation, indicating that the phonon splitting mechanisms for the $E_{g}$ and $E_{u}$ modes are distinct.
We note that the energy splitting of the $E_{u}$ mode is significantly smaller than that of the $E_{g}$ mode at $\Gamma$ point, a result captured by both our ab initio calculations and experimental measurements [8, 9]. Moreover, We notice both the $E_{g}$ and $E_{u}$ modes exhibit a broad distribution of splitting across the entire BZ, rather than being concentrated solely at the high-symmetry momentum. This widely distributed chiral phonon dynamics may give rise to observable transport signatures associated with phonon chirality.
Fig. 4(c) displays the distributions of pseudo-angular momentum and angular momentum for the two split $E_{u}$ modes. The angular momentum of these modes is not quantized to unity, which can be attributed to the vibration from both Co and S atoms. As illustrated in Fig. 4(d), the Co atoms are not located at $C_{3}$-invariant Wyckoff positions, thus resulting in a non-quantized values of angular momentum [23].
Figure 5: Distribution of the MBC in the electronic BZ. (a-c) are MBC distributions for the $E_{g}$ phonon mode with $l_{ph}=+1$ on three distinct $k_{z}$ planes. The MBC distribution for the $E_{g}$ mode with $l_{ph}=-1$ is of opposite sign. (d-f) are MBC distributions for the $E_{u}$ mode with $l_{ph}=+1$. The $E_{u}$ mode with $l_{ph}=-1$ also exhibits a sign reversal. The MBC magnitude is significantly larger for the $E_{g}$ modes than for the $E_{u}$ modes, which consequently results in a smaller calculated phonon splitting for the $E_{u}$ modes.
Phonon splitting and magnetism under MBC To explain the distinct magnitude of the energy splitting for the $E_{g}$ and $E_{u}$ modes at the $\Gamma$ point, we calculate the distribution of the mode-resolved MBC, denoted as $\tilde{G}(\Gamma,\bm{k},\bm{k})$, across the electronic BZ, as shown in Fig. 5. The results indicate that $\tilde{G}_{E_{u}}(\Gamma,\bm{k},\bm{k})$ is significantly smaller in magnitude than $\tilde{G}_{E_{g}}(\Gamma,\bm{k},\bm{k})$, consistent with our earlier findings on MBC-induced phonon splitting. Furthermore, we observe that the distribution of $\tilde{G}_{E_{u}}(\Gamma,\bm{k},\bm{k})$ exhibits substantial cancellation between positive and negative contributions. This cancellation further explains the markedly smaller MBC-induced splitting of the $E_{u}$ mode compared to the $E_{g}$ mode.
Previous studies have attributed $\Gamma$-point phonon splitting under external magnetic fields to the manifestation of an effective phonon magnetic moment ($M_{z}$[3, 4, 1, 2]. As detailed in Ref. [3], $M_{z}$ comprises two contributions: one from the Born effective charges and another from the Berry curvature of electrons in both real space (MBC) and momentum space ($\Omega_{e}$), i.e.,
$M_{z}=\frac{e}{2M_{\kappa}}l_{z,\nu\bm{q}}\int d\bm{k}\left\{\Omega_{k_{x}u_{y}}-\Omega_{k_{y}u_{x}}+\Omega_{e,k_{x}k_{y}}G_{u_{x}u_{y}}\right\}.$ (4)
The first two terms collectively integrate to the Born effective charge, whereas the third term is equivalent to the product of the conventional electronic and molecular Berry curvatures. In css, the metallic nature and the preservation of $\mathcal{P}$ result in a vanishing Born effective charge. Given that $G$ is independent of $\bm{k}$ and the integral $\int d\bm{k}\Omega_{e,k_{x}k_{y}}$ is proportional to the anomalous Hall conductivity $\sigma^{\text{A}}_{xy}$, the expression for the phonon magnetic moment simplifies to $M_{z}\propto l_{z,\nu\bm{q}}\cdot G\cdot\sigma^{\text{A}}_{xy}$. Consequently, the large intrinsic Hall conductivity of css (approximately 505 $\Omega^{-1}$ cm-1 [25]) indicates a strong propensity for intrinsic phonon magnetism in this material. As calculated and shown in Fig. 5, the pronounced difference in the MBC between the $E_{g}$ and $E_{u}$ modes indicates a larger magnetic moment for the $E_{g}$ phonon.
In summary, our ab initio-MBC methodology captures the electronic order-induced symmetry breaking in the lattice dynamics of css, generalizing the manifestation of phonon magnetism and phonon energy shift effects in magnetic materials. Notably, the energy scale of circularly polarized phonons coincides with that of chiral Weyl fermions, suggesting rich opportunities for chiral boson–fermion coupling and enabling control over topological states through phononic and magnetic means [47]. This efficient and accurate framework for magnetic lattice dynamics is readily extendable to full Brillouin-zone calculations and enables the study of emergent phenomena such as the phonon Hall effect [13, 14, 15, 16, 17, 15, 18, 19, 20, 4], Hall viscosity [48], and so on. The algorithm is broadly applicable to intrinsic magnetic systems, facilitating the discovery of materials with large phonon magnetism.
Our findings establish a set of design principles for materials with large phonon magnetic moments and substantial splitting, i.e., strong spin-orbit coupling, significant electron-phonon coupling, and phonon modes carrying nonzero angular momentum. We also propose that materials integrating magnetic order with flat bands, such as css, represent a particularly fertile platform for the discovery of giant phonon splitting and pronounced phonon magnetism.
Future studies may focus on extrinsic contributions to the phonon magnetic moment, such as relaxation processes [4], non-reciprocal transport related to circularly polarized phonons [49, 21, 50, 51], and ab initio algorithm for external field–induced phonon magnetism, etc.
I Acknowledgement
We acknowledge the helpful discussion with Haijun Zhang, Fuyi Wang, Xi Dai, Yang Gao, Yafei Ren, Shang Ren, Run Yang and Martin Dressel. We acknowledge the support from National Key R&D Project (Grant Nos. 2023YFA1407400 and 2024YFA1409200), and the National Natural Science Foundation of China Grant No. 12374165.

### v3

@@
-Electronic order induced symmetry breaking in lattice dynamics of Co3Sn2S2
+General ab initio framework for chiral phonons induced by electronic order
 Abstract
-Based on the molecular Berry curvature (MBC) framework, we develop an ab initio algorithm to capture the quantitative effects of magnetic order on lattice dynamics.
-Using the ferromagnetic Weyl semimetal Co3Sn2S2 as a prototype, we show that electronic-order-driven phonon symmetry breaking requires spin-orbit coupling (SOC) and leads to an MBC term that breaks both time-reversal ($\mathcal{T}$) and mirror symmetries.
-We demonstrate that mirror-symmetry breaking is essential to account for the experimentally observed phonon splitting, $\mathcal{T}$-breaking alone is insufficient.
-The MBC is widely distributed across the Brillouin zone, giving rise to significant off-$\Gamma$ effects.
-Our results show the phonon splitting of the $E_{g}$ and $E_{u}$ modes have distinct origins: the $E_{g}$ splitting arises from MBC and is captured by our algorithm, while the $E_{u}$ splitting stems from correlation-sensitive Fano resonance, explaining their asymmetric experimental line shapes.
-We provide a design strategy and computational pathway for materials with large phonon magnetism and splitting, requiring SOC, electron-phonon coupling, and ideally flat bands at the Fermi level.
-This work also suggests new avenues for controlling non-reciprocal phonon transport.
+Conventional ab initio methods fail to describe the emergence of chiral phonons driven by electronic ordering.
+Here, we develop an ab initio framework, grounded in molecular Berry curvature (MBC), that captures electronic-order-driven symmetry breaking in lattice dynamics and is applicable to both insulating and metallic magnets.
+Using Co3Sn2S2 as a model system, we show that the MBC term simultaneously breaks time-reversal and mirror symmetries, enabling a quantitative reproduction of the experimentally observed phonon splittings.
+The analysis uncovers distinct microscopic origins for the $E_{g}$ and $E_{u}$ modes: the $E_{g}$ splitting is governed by MBC and is accurately described by our first-principles scheme, whereas the $E_{u}$ splitting is enhanced by the Fano resonance, consistent with its asymmetric spectral profile.
+Leveraging this framework, we further predict several candidates with chiral-phonon splitting.
+Our results establish a predictive route for identifying and understanding phonon magnetism, chiral phonons, and related Hall-type lattice responses from first principles.
+I Introduction
 Within Landau’s theory, electronic orders, such as magnetic/spin, charge, and orbital order, are typically characterized by local order parameters and strong electron interactions, and are invariably accompanied by spontaneous symmetry breaking.
 This contrasts with phonon systems, where such symmetry breaking is rare.
-Recent work has shown that time-reversal symmetry ($\mathcal{T}$) breaking occurs in phonons by coupling to external fields [1, 2, 3, 4], spins [5, 6] and chiral excitations [7].
-Moreover, phonon energy splitting in ferromagnetic (FM) Weyl semimetal css directly demonstrates this effect [8, 9].
-Although ab initio methods can readily describe electronic symmetry breaking, such as $\mathcal{T}$ in magnetic systems, the concomitant symmetry breaking in the phonon system induced by electronic order remains inadequately explored by first-principles calculations.
-Thus, a quantitative understanding of symmetry breaking in lattice dynamics from first principles [10, 11, 12] is essential, and it is crucial for interpreting phonon-related transport [13, 14, 15, 16, 17, 15, 18, 19, 20, 4] and optical properties [9, 8, 21, 22, 1].
-To capture the experimentally observed phonon splitting in css, which is induced by magnetic order, we develop an ab initio algorithm based on molecular Berry curvature (MBC, or nuclear Berry curvature) theory.
-A nonzero MBC contribution necessitates the inclusion of spin-orbit coupling (SOC) [10].
-This requirement distinguishes our approach from the phonon self-energy method in Ref.
-[11], which yields phonon splitting even in the absence of SOC.
-Under FM order and SOC, the MBC contribution breaks both $\mathcal{T}$ and mirror symmetries in css.
-Our results demonstrate that this mirror-symmetry breaking is essential to explain the observed phonon splitting in css, as $\mathcal{T}$ breaking alone is insufficient [23].
-Furthermore, there’s a broad Brillouin zone (BZ) distribution of the MBC contributions, leading to significant effects away from the $\Gamma$-point.
-Both mirror-symmetry breaking and off-$\Gamma$ effects, which remain largely unexplored, are critical for a complete theoretical description of the lattice dynamics in css.
-Furthermore, our results demonstrate that the magnitude of the phonon splitting, while relevant for circularly polarized phonons with nonzero angular momentum (i.e., chiral phonons), is not uniquely determined by its absolute value.
-Figure 1: The crystal structure, Brillouin zone, electronic/phonon structure of css.
-(a) Crystal structure of css.
-Brown arrows indicate the local magnetic moments on Co atoms.
-This magnetic order breaks the vertical mirror and time-reversal symmetry, but keeps the $C_{3}$ and inversion symmetry.
-(b) Brillouin zone of css.
-(c) Electronic band structure with FM order and spin-orbit coupling.
-(d) Phonon spectrum of css without the molecular Berry curvature contribution.
-The coordinates of $K$ are ($\frac{1}{3}$, 0, -$\frac{1}{3}$)/ ($\frac{1}{3}$, $\frac{1}{3}$, 0) in the basis of the primitive/conventional lattice vectors.
-Lattice dynamics for css without MBC css is the first experimentally confirmed magnetic Weyl semimetal [24, 25, 26, 27, 28, 29, 30, 31, 32, 33] with a centrosymmetric space group of $R\bar{3}m$ ($\mathcal{P}$-preserving), thus the phonon modes can be classified in Raman active ones and infrared (IR) active ones at $\Gamma$.
+Recent work has shown that time-reversal symmetry ($\mathcal{T}$) breaking occurs in phonons by coupling to external fields [Baydin2022_PRB, Cheng2020_Cd3As2, Luo2023_CeF3_Science, Chaudhary2024_PRB, Wu2023_Fe2Mo3O8_NP, David2024_CoTiO3_PNAS, Nova2017_ErFeO3_NP, Schaack1977_CeCl3, CeCl3_2022, Niu2021_phononMag, Xue2025_Extrinsic], spins [Ren2024_PH_SPIN_PRX, Liu2021_Magnon_phonon] and chiral excitations [FeGeTe_2D_chiral_phonon_Du_2019].
+Moreover, phonon energy splitting in ferromagnetic (FM) Weyl semimetal css directly demonstrates this effect [CoSnS_CP_2025, CoSnS_PRL_2025].
+Although ab initio methods can readily describe electronic symmetry breaking, such as $\mathcal{T}$ in magnetic systems, the concomitant symmetry breaking in the phonon system induced by electronic order remains inadequately explored.
+Thus, a quantitative understanding of symmetry breaking in lattice dynamics from first principles [JunRen_Shi2017, Bistoni2021_Noncollinear_PRL, Bonini2023_CrI3_TRS, Ren2024_PH_SPIN_PRX, HJZhang2025_EPC, chen2025magicnonlocalgeometricforce] is essential, and it is crucial for interpreting phonon-related transport [Strohm2005_THE, Nagaosa2010_THE_QuantumMagnets, Ideue2017_THE_multiferroics, Zhang2021_ATHE_VI3, Nagaosa2019_Berry_THE, Niu2022_MolecularBerry, Xue2025_Extrinsic] and optical properties [CoSnS_CP_2025, CoSnS_PRL_2025, nagaosa2024nonreciprocal, CeCl3_2022].
+In this work, we develop an ab initio algorithm based on molecular/nuclear Berry curvature (MBC) theory [MT1979_lattice_dyn, Niu2022_MolecularBerry] to capture the magnetism-induced symmetry breaking in lattice dynamics, which necessitates the inclusion of spin-orbit coupling (SOC) [Ren2024_PH_SPIN_PRX].
+Using Co3Sn2S2 as a representative system, we show that the MBC contribution breaks both $\mathcal{T}$ and mirror symmetries under FM order and SOC, and that mirror-symmetry breaking is essential to reproduce the observed phonon splitting, as $\mathcal{T}$ breaking alone is insufficient [Sinisa_PT_AM_2023, Zhang2025_CP].
+The phonon splitting extends broadly across the Brillouin zone, and together with MBC-induced mirror-symmetry breaking, is essential for a complete theoretical description of lattice dynamics in Co3Sn2S2.
+This approach further predicts candidate materials exhibiting circularly polarized phonon splitting driven by magnetic order, SOC, and electron–phonon coupling.
+Our results show that although phonon splitting requires a mode with nonzero angular momentum, the magnitude of the splitting is not determined by the angular momentum alone.
+Instead, the splitting amplitude arises from the combined effects of magnetic order, SOC, and electron–phonon interaction.
+II Results
+II.1 Lattice dynamics without MBC
+In standard first-principles calculations, the dynamical matrix $K_{ij}$ is constructed to preserve all crystal symmetries and $\mathcal{T}$, even when these are broken by FM order.
+This algorithmic limitation leads to inaccurate phonon spectra.
+For instance, the lattice model shown in Fig.
+1(a) possesses $C_{3}$ rotational, inversion ($\mathcal{P}$), vertical mirror ($\sigma_{v}$), and $\mathcal{T}$ symmetries.
+Ferromagnetic order along the out-of-plane direction breaks both $\sigma_{v}$ and $\mathcal{T}$.
+However, conventional first-principles approaches still enforce these symmetries in the dynamical matrix, as in the nonmagnetic case, yielding the phonon spectra shown in Fig.
+1(b), which fail to capture the symmetry breaking intrinsic to the FM state.
+The accurate phonon spectra, reflecting the true lattice dynamics of the magnetic system, should correspond to Fig.
+1(c), which lies beyond the capability of standard first-principles methods.
+In the following, we illustrate how the MBC theory corrects lattice dynamics and apply it to Co3Sn2S2 as a representative example, demonstrating that our numerical calculations capture the mechanisms underlying the experimentally observed phonon splittings.
+Figure 1: Schematic illustration of lattice dynamics in magnetic materials under the standard ab initio approach and the MBC-corrected framework.
+(a) The model lattice preserves $C_{3}$, inversion ($\mathcal{P}$), vertical mirror ($\sigma_{v}$), and time-reversal ($\mathcal{T}$) symmetries.
+Ferromagnetic order along the out-of-plane direction breaks both $\sigma_{v}$ and $\mathcal{T}$.
+(b) Phonon spectra obtained by standard ab initio calculations, where the doubly degenerate phonon dispersion is enforced by $\sigma_{v}$ and $\mathcal{T}$ in the dynamical matrix $K(\bm{q})$.
+(c) By introducing the MBC-related term $G(\bm{q})$ into the lattice dynamics, our algorithm accurately captures the phonon spectra of magnetic materials.
+The MBC term generates opposite phonon magnetic moments $M_{ph}$ for right- and left-handed rotational vibrations, thereby breaking $\sigma_{v}$ and $\mathcal{T}$ symmetries.
+(d) Schematic illustration of the molecular Berry curvature arising from the adiabatic evolution of the electronic ground state $\Phi_{0}$ under different phonon modes.
+Phonon modes with opposite circular polarizations acquire MBC corrections of opposite sign.
+css is the first experimentally confirmed magnetic Weyl semimetal [Xu2018_PRB_CoSnS_Weyl, Liu2018, Wang2018_NC, science.aav2873, science.aav2334, YAN201857, PhysRevLett.124.077403, PhysRevMaterials.4.044203, Xu2020, Okamura2020_magopt_CSS] with a centrosymmetric space group of $R\bar{3}m$ ($\mathcal{P}$-preserving), thus the phonon modes can be classified in Raman active ones and infrared (IR) active ones by the $\mathcal{P}$ parity at $\Gamma$.
 Figure 2(a) shows the primitive cell of css with brown arrows indicating the local magnetic moment.
-Below $T_{\rm C}\sim 175$[34, 27, 28, 29, 35, 36, 37, 38], FM order with $C_{3}$ axis spin alignment breaks both $\mathcal{T}$ and vertical mirror symmetry, enabling an SOC-essential Weyl semimetal phase, reducing the point group from $D_{3d}$ to $C_{3i}$ (magnetic space group $R\bar{3}m^{\prime}$).
-Figure 1(c) presents the electronic band structure of css with FM order and SOC.
-The Fermi level has been shifted by 25 meV to align with the in-plane flat band position observed by the scanning tunneling microscopy [39].
-The electronic bands near the Fermi level are primarily contributed by the $d$-orbitals of Co atoms.
-Figure 1(d) displays the phonon spectrum without the MBC contribution.
-The highlighted $E_{u}$ and $E_{g}$ modes correspond to those observed experimentally [9, 8], and will be discussed in detail through quantitative comparisons.
-The sole $E_{g}$ mode arises from S atoms occupying the $6c$ Wyckoff positions, which lacks inversion site-symmetry.
-In contrast, Co, Sn1 and Sn2 reside at the Wyckoff positions of $9e$, $3a$ and $3b$, respectively, all of them preserve inversion site-symmetry.
-The atomic displacements transform as a vector representation under $O(3)$ operations, analogous to $p$-orbitals, and exhibit odd parity.
-As a result, neither Co nor Sn can generate phonon modes of even parity.
-The degeneracy of the $E_{g}$ and $E_{u}$ modes along the $C_{3}$-invariant path, i.e., $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$, is protected by vertical mirror symmetry as well as $\mathcal{PT}$ symmetry [23].
-The experimental observation of phonon splitting thus implies the breaking of both $\mathcal{PT}$ and vertical mirror symmetry, as shown in Fig.
-2(e).
-To quantitatively describe the underlying physical mechanism, we develop a ab initio+MBC algorithm based on the Wannier basis to efficiently quantify this ferromagnetism-induced $E_{g/u}$ splitting.
-Figure 2: Phonon spectra near the $\Gamma$ point under successive symmetry-breaking conditions.
-From (a) to (c), vertical mirror symmetry ($\sigma_{v}$) and inversion symmetry ($\mathcal{P}$) are progressively broken.
-From top to bottom, time-reversal symmetry ($\mathcal{T}$) is broken.
-Since css belongs to the $D_{3d}$ point group, breaking $\mathcal{T}$ alone is insufficient to lift the phonon degeneracy, as seen between (a) and (d).
-Therefore, the MBC mechanism, which involves spin–orbit coupling and breaks $\sigma_{v}$, is necessary to split the degenerate modes.
-MBC modified lattice dynamics Experiments rule out resonant magnon-phonon coupling in css [40], allowing it to be safely discounted for an accurate description of the lattice dynamics [10, 2].
-Furthermore, Ref.
-[9] confirms weak electron-hole pair excitation in Co3Sn2S2, preserving Born-Oppenheimer approximation (BOA) validity for MBC calculations.
-Within the BOA, the adiabatic evolution of the electronic ground state with the lattice accumulates a Berry phase, which subsequently couples back to influence lattice dynamics [41, 20], thus the equation of motion (EOM) reads:
-$\tilde{K}(\bm{q})+\tilde{G^{\dagger}}(\bm{q})\tilde{G}(\bm{q})]\epsilon_{\nu\bm{q}}=\omega^{2}_{\nu\bm{q}}\epsilon_{\bm{q}}-2i\omega_{\bm{q}}\tilde{{G}}(\bm{q})\epsilon_{\nu\bm{q}}$}, (1)
+Below $T_{\rm C}\sim 175$[Liu2018], FM order with $C_{3}$ axis spin alignment breaks both $\mathcal{T}$ and vertical mirror symmetry, enabling an SOC-essential Weyl semimetal phase, reducing the point group from $D_{3d}$ to $C_{3i}$ (magnetic space group $R\bar{3}m^{\prime}$).
+Figure 2(c) shows the phonon spectrum of ferromagnetic Co3Sn2S2 with SOC, obtained from standard ab initio calculations, which remains degenerate as in Fig.
+1(b).
+The highlighted $E_{u}$ and $E_{g}$ modes correspond to those observed experimentally [CoSnS_CP_2025, CoSnS_PRL_2025]; however, the experimental phonon modes are non-degenerate, analogous to Fig.
+1(c).
+The degeneracy of the Eg/uE{g/u} modes along the $C_{3}$-invariant path ($\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$) is protected by vertical mirror and $\mathcal{PT}$ symmetries [Zhang2025_CP].
+Therefore, the experimentally observed phonon splitting indicates the breaking of both symmetries, as illustrated in Fig.
+1(c).
+II.2 MBC modified lattice dynamics
+To capture lattice dynamics in magnets accurately, the Berry phase arising from adiabatic electronic evolution must be included [EM_THE_Qin2011_PRL, Niu2022_MolecularBerry], as it feeds back into the lattice motion and yields the $G(\bm{q})$ term shown in Figs.
+1(c) and (d).
+Thus the equation of motion (EOM) reads:
+$\tilde{K}(\bm{q})+{2}\tilde{G^{\dagger}}(\bm{q})\tilde{G}(\bm{q}){+2i\omega_{\nu\bm{q}}\tilde{{G}}(\bm{q})}]\epsilon_{\nu\bm{q}}=\omega^{2}_{\nu\bm{q}}\epsilon_{\bm{q}},$ (1)
 where $\epsilon_{\nu\bm{q}}$ is the polarization vector, $\tilde{K}_{\kappa\alpha,\kappa^{\prime}\beta}(\bm{q})=\frac{1}{\sqrt{M_{\kappa}M_{\kappa^{\prime}}}}K_{\kappa\alpha,\kappa^{\prime}\beta}(\bm{q})$ and $\tilde{G}_{\kappa\alpha,\kappa^{\prime}\beta}(\bm{q})=\frac{1}{2\sqrt{M_{\kappa}M_{\kappa^{\prime}}}}G_{\kappa\alpha,\kappa^{\prime}\beta}(\bm{q})$ are the mass-weighted dynamic matrix and MBC.
@@
 $\nu=\nu^{\prime}$ for the non-degenerated case. Detailed discussions are in the Supplementary Material .
-Standard first-principles calculations of the dynamical matrix $K_{ij}$ preserve the symmetries that are broken by the electronic FM order parameterized via a pseudovector, such as $\mathcal{T}$ and vertical mirror symmetry (see Supplementary Material).
-Thus, this algorithmic limitation results in a spurious double degeneracy of the $E_{g}$ and $E_{u}$ phonon modes, as shown in Fig.
-2(a).
-Although the MBC term $\tilde{G}$ was proposed to break $\mathcal{T}$-symmetry [20, 20, 10], it alone is insufficient to lift the degeneracy of phonon bands along the $C_{3v}$-invariant high-symmetry path ($\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$) in css, due to the presence of vertical mirror symmetry [23], as illustrated in Fig.
-2 (d).
-Therefore, an additional mechanism breaking the both of the $\mathcal{T}$ and vertical mirror symmetry by MBC is essential to explain the experimentally observed phonon splitting in Fig.
-2 (e).
-Figure 3: The MBC induced splitting for $E_{g}$ mode.
-(a) Phonon dispersion of the $E_{g}$ mode along the $C_{3}$-invariant path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$, calculated without (black dashed line) and with molecular Berry curvature (blue and red solid lines).
+Figure 2: (a) Crystal structure of css.
+Brown arrows indicate the local magnetic moments on Co atoms.
+This magnetic order breaks the vertical mirror and time-reversal symmetry, but keeps the $C_{3}$ and inversion symmetry.
+(b) Brillouin zone of css.
+(c) Phonon spectrum of FM-ordered css with SOC, excluding the molecular Berry curvature contribution.
+The coordinate of $K$ is ($\frac{1}{3}$, 0, -$\frac{1}{3}$)/ ($\frac{1}{3}$, $\frac{1}{3}$, 0) in the basis of the primitive/conventional lattice vectors.
+(d1) Phonon dispersion of the $E_{g}$ mode along the $C_{3}$-invariant path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$, calculated without (black dashed line) and with molecular Berry curvature (blue and red solid lines).
 The colored stars mark the energy positions of the $E_{g}$ mode observed experimentally.
-(b) Energy splitting of the $E_{g}$ mode along the $C_{3}$-invariant path.
+(d2) Energy splitting of the $E_{g}$ mode along the $C_{3}$-invariant path.
 The insert figure shows the vibrational modes of the S atoms at the $\Gamma$ and T points.
-(c) Evolution of the pseudo-angular momentum ($l_{ph}$) and the $z$-component of the angular momentum ($l_{z,\nu\bm{q}}$) along the $C_{3}$-invariant path.
-(d) Atomic displacement patterns of the two $E_{g}$ modes at the $\Gamma$ point.
-MBC-modified $E_{g}$ mode Figure 3(a) displays the phonon spectra of the $E_{g}$ mode with MBC along the $C_{3}$-invariant path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$.
-Within our ab initio framework, MBC lifts the degeneracy along the entire path by breaking both time-reversal $\mathcal{T}$ (as well as $\mathcal{PT}$) and vertical mirror symmetries.
+(d3) Evolution of the pseudo-angular momentum ($l_{ph}$) and the $z$-component of the angular momentum ($l_{z,\nu\bm{q}}$) along the $C_{3}$-invariant path.
+(d4) Atomic displacement patterns of the two $E_{g}$ modes at the $\Gamma$ point.
+(e1)-(e4) are the corresponding contents for the $E_{u}$ mode.
+Experiments rule out resonant magnon-phonon coupling in css [Liu2020spin], allowing it to be safely discounted for an accurate description of the FM-ordered lattice dynamics [Ren2024_PH_SPIN_PRX, Bonini2023_CrI3_TRS].
+Furthermore, Ref.
+[CoSnS_CP_2025] confirms weak electron-hole pair excitation in Co3Sn2S2, preserving Born-Oppenheimer approximation (BOA) validity for MBC calculations.
+Although the MBC term $\tilde{G}$ was proposed to break $\mathcal{T}$-symmetry [Niu2022_MolecularBerry, Niu2022_MolecularBerry, Ren2024_PH_SPIN_PRX], it alone is insufficient to lift the degeneracy of phonon bands along $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$ due to the presence of $\sigma_{v}$ [Zhang2025_CP].
+Therefore, the MBC-induced breaking of $\sigma_{v}$ is essential for explaining the phonon splitting in experiments.
+II.2.1 MBC-modified $E_{g}$ mode
+Figure 2(d1) displays the phonon spectra of the $E_{g}$ mode with MBC along $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$.
+Within our ab initio framework, MBC lifts the degeneracy along the entire path by breaking both $\mathcal{T}$ (as well as $\mathcal{PT}$) and $\sigma_{v}$.
 The magnitude of the phonon energy splitting for the $E_{g}$ modes is presented in Fig.
-3(b), revealing a maximum splitting at the $\Gamma$ point and a minimum splitting at T.
+2(d2), revealing a maximum splitting at the $\Gamma$ point and a minimum splitting at T.
 At $\Gamma$, the calculated phonon splitting is approximately 0.253 meV (2.05 cm-1), which agrees with the experimental value of 1.27 cm-1 reported in Ref.
-[8] in order of magnitude.
-The minor discrepancy may originate from slight variations in the experimental Fermi level and anharmonic effects [42, 43].
-To further investigate the influence of the Fermi level on the MBC, we shifted it by 60 meV and recalculated the splitting of the $E_{g}$ mode.
-This adjustment yields a splitting of approximately 1.24 cm-1, which is in excellent agreement with the experimental value (see Supplementary Material for details).
+[CoSnS_PRL_2025] in order of magnitude.
+The minor discrepancy may originate from slight variations in the experimental Fermi level and anharmonic effects.
+To further investigate the influence of the Fermi level on the MBC, we shift it by 60 meV and recalculate the splitting of the $E_{g}$ mode.
+This adjustment yields a splitting of approximately 1.24 cm-1, which is in excellent agreement with the experimental value (see Supplementary Material Fig.
+LABEL:fig:Eg_ef60 for details).
 To further elucidate the lattice dynamics of $E_{g}$ induced by MBC, we calculate both the pseudo-angular momentum ($l_{ph}$) and the $z$-component of the angular momentum ($l_{z,\nu\bm{q}}$) along the $C_{3}$-invariant path, as shown in Fig.
-3(c).
-We note the angular momentum only has $z$ component due to the restriction of $C_{3}$ symmetry [23].
-Both quantities exhibit a quantized value of unity, with sign reversals resulting from band inversion along the path, which fundamentally governs the emergence of the phonon Hall effect [13, 14, 15, 16, 17, 15, 18, 19, 20, 4].
-The quantized unity of the $l_{\text{ph}}$ arises from the $C_{3}$ symmetry of the system.
-In contrast, the quantization of $l_{z,\nu\bm{q}}$ along the path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$ can be attributed to two factors: (i) At $\Gamma$, $E_{g}$ is contributed solely by the vibration of S atoms residing at $C_{3}$-invariant Wyckoff positions [44, 23].
+2(d3).
+We note the angular momentum only has $z$ component due to the restriction of $C_{3}$ symmetry [Zhang2025_CP].
+Both quantities exhibit a quantized value of unity, with sign reversals resulting from band inversion along the path, which fundamentally governs the emergence of the Weyl phonon [zhang2018double, FeSi_hmiao, zhang2025new] and phonon Hall effect [Strohm2005_THE, Nagaosa2010_THE_QuantumMagnets, Tao2012_Berry_THE, Ideue2017_THE_multiferroics, Zhang2021_ATHE_VI3, Nagaosa2019_Berry_THE, Niu2022_MolecularBerry, Xue2025_Extrinsic].
+The quantized unity of the $l_{\text{ph}}$ arises from the $C_{3}$ symmetry of the system, while the quantization of $l_{z,\nu\bm{q}}$ along the path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$ can be attributed to two factors: (i) At $\Gamma$, $E_{g}$ is contributed solely by the vibration of S atoms residing at $C_{3}$-invariant Wyckoff positions [zhang2022chiral, Zhang2025_CP].
 As illustrated in Fig.
-3(d), these two S atoms rotate in the same direction but with a $\pi$ phase difference at the $\Gamma$.
+2(d4), these two S atoms rotate in the same direction but with a $\pi$ phase difference at the $\Gamma$.
 (ii) Away from $\Gamma$, their energies are well-separated from other modes, resulting in minimal hybridization with other phonon modes.
@@
 As shown in Figs.
-3(b) and (c), $l_{z,\nu\bm{q}}$ remains close to unity along the $C_{3}$-invariant path, yet the splitting of the $E_{g}$ modes varies significantly with a maximum at $\Gamma$.
-Figure 4: The MBC induced splitting for $E_{u}$ mode.
-(a) Phonon dispersion of the $E_{u}$ mode along the $C_{3}$-invariant path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$, calculated without (black dashed line) and with molecular Berry curvature (blue and red solid lines).
-The colored stars mark the energy positions of the $E_{u}$ mode observed experimentally.
-(b) Energy splitting of the $E_{u}$ mode along the $C_{3}$-invariant path.
-The insert figure shows the vibrational modes of the S atoms at the $\Gamma$ and T points.
-(c) Evolution of the pseudo-angular momentum ($l_{\text{ph}}$) and the $z$-component of the angular momentum ($l_{z,\nu\bm{q}}$) along the $C_{3}$-invariant path.
-(d) Atomic displacement pattern of the two $E_{u}$ modes at the $\Gamma$ point.
-MBC-modified $E_{u}$ mode Figure 4(a) displays the MBC-modified phonon spectra of the $E_{u}$ mode along the path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$.
-The MBC also lifts the degeneracy of the $E_{u}$ mode (dashed line) across the entire $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$ path.
-The magnitude of the phonon energy splitting for the Eg modes is presented in Fig.
-4(b), revealing a maximum splitting in the vicinity of T and minimum splitting in the vicinity of $\Gamma$.
-Combined with Figs.
-3(b) and  4(b), it can be observed that although the parity of the $E_{g}$ and $E_{u}$ modes remains unchanged at both $\Gamma$ and T, the relative phase between the two S atoms shifts by $\pi$.
-For instance, the $E_{g}$ mode exhibits a relative phase of $\pi$ at $\Gamma$ but 0 at T, whereas the $E_{u}$ mode shows a phase of 0 at $\Gamma$ and $\pi$ at T.
-This switching of the relative phase between the two modes correlates with the reversal of the maximum phonon splitting.
-At the $\Gamma$ point, the splitting calculated from our algorithm is about 0.002 meV (0.016 cm-1), which is smaller than the experimentally observed value of 0.3 cm-1 reported in the optical conductivity measurements [9].
+2(d2) and (e2), $l_{z,\nu\bm{q}}$ remains close to unity along the $C_{3}$-invariant path, yet the splitting of the $E_{g}$ modes varies significantly with a maximum at $\Gamma$.
+II.2.2 MBC-modified $E_{u}$ mode
+Figure 2(e1) displays the MBC-modified phonon spectra of the $E_{u}$ mode along the path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$, where the MBC also lifts the degeneracy (dashed line) across the entire path.
+The magnitude of the phonon energy splitting for the $E_{u}$ modes is presented in Fig.
+2(e2), revealing a maximum splitting in the vicinity of T and minimum splitting in the vicinity of $\Gamma$.
+Both $E_{g/u}$ modes exhibit a broad distribution of splitting across the entire BZ, rather than being concentrated solely at the high-symmetry momentum.
+This widely distributed chiral phonon dynamics may give rise to observable transport signatures associated with phonon chirality.
+At the $\Gamma$ point, the splitting calculated from our algorithm is about 0.002 meV (0.016 cm-1), which is smaller than the experimentally observed value of 0.3 cm-1 reported in the optical conductivity measurements [CoSnS_CP_2025].
 This discrepancy suggests that the splitting observed in experiments is not primarily caused by the MBC.
-Combined with the asymmetric line shapes of the $E_{u}$ mode, we attribute the larger experimental splitting to a Fano resonance [9].
-This resonance may arise from quantum interference between electronic and phonon excitations when Weyl cones approach the Fermi level [45, 46].
+Combined with the asymmetric line shapes in experiments, we attribute the larger experimental splitting to the Fano resonance, which may arise from quantum interference between electronic and phonon excitations when Weyl cones approach the Fermi level [cappelluti_charged2012, CoSnS_CP_2025].
 The absence of an asymmetric line shape in the $E_{g}$ mode further supports this interpretation, indicating that the phonon splitting mechanisms for the $E_{g}$ and $E_{u}$ modes are distinct.
-We note that the energy splitting of the $E_{u}$ mode is significantly smaller than that of the $E_{g}$ mode at $\Gamma$ point, a result captured by both our ab initio calculations and experimental measurements [8, 9].
-Moreover, We notice both the $E_{g}$ and $E_{u}$ modes exhibit a broad distribution of splitting across the entire BZ, rather than being concentrated solely at the high-symmetry momentum.
-This widely distributed chiral phonon dynamics may give rise to observable transport signatures associated with phonon chirality.
-Fig.
-4(c) displays the distributions of pseudo-angular momentum and angular momentum for the two split $E_{u}$ modes.
-The angular momentum of these modes is not quantized to unity, which can be attributed to the vibration from both Co and S atoms.
-As illustrated in Fig.
-4(d), the Co atoms are not located at $C_{3}$-invariant Wyckoff positions, thus resulting in a non-quantized values of angular momentum [23].
+Figure 2(e3) displays the distributions of $l_{ph}$ and $l_{z,\nu\bm{q}}$ for the two split $E_{u}$ modes, where $l_{z,\nu\bm{q}}$ is not quantized to unity.
+This can be attributed to the vibration from both Co and S atoms, and Co atoms are not located at $C_{3}$-invariant Wyckoff positions [Zhang2025_CP], as illustrated in Fig.
+2(e4).
 Figure 3: Distribution of the MBC in the electronic BZ.
@@
 The $E_{u}$ mode with $l_{ph}=-1$ also exhibits a sign reversal.
-The MBC magnitude is significantly larger for the $E_{g}$ modes than for the $E_{u}$ modes, which consequently results in a smaller calculated phonon splitting for the $E_{u}$ modes.
-Phonon splitting and magnetism under MBC To explain the distinct magnitude of the energy splitting for the $E_{g}$ and $E_{u}$ modes at the $\Gamma$ point, we calculate the distribution of the mode-resolved MBC, denoted as $\tilde{G}(\Gamma,\bm{k},\bm{k})$, across the electronic BZ, as shown in Fig.
-5.
+The MBC magnitude is significantly smaller for the $E_{u}$ modes, which consequently results in a smaller calculated phonon splitting.
+II.3 Phonon splitting and magnetism from MBC
+For css, the energy splitting of the $E_{u}$ mode is significantly smaller than that of the $E_{g}$ mode at $\Gamma$ point, a result captured by both our ab initio calculations and experimental measurements [CoSnS_PRL_2025, CoSnS_CP_2025].
+To explain their distinct magnitude of the energy splitting, we calculate the distribution of the mode-resolved MBC, denoted as $\tilde{G}(\Gamma,\bm{k},\bm{k})$, across the electronic BZ, as shown in Fig.
+3.
 The results indicate that $\tilde{G}_{E_{u}}(\Gamma,\bm{k},\bm{k})$ is significantly smaller in magnitude than $\tilde{G}_{E_{g}}(\Gamma,\bm{k},\bm{k})$, consistent with our earlier findings on MBC-induced phonon splitting.
-Furthermore, we observe that the distribution of $\tilde{G}_{E_{u}}(\Gamma,\bm{k},\bm{k})$ exhibits substantial cancellation between positive and negative contributions.
-This cancellation further explains the markedly smaller MBC-induced splitting of the $E_{u}$ mode compared to the $E_{g}$ mode.
-Previous studies have attributed $\Gamma$-point phonon splitting under external magnetic fields to the manifestation of an effective phonon magnetic moment ($M_{z}$[3, 4, 1, 2].
+Furthermore, we observe that the distribution of $\tilde{G}_{E_{u}}(\Gamma,\bm{k},\bm{k})$ exhibits substantial cancellation between positive and negative contributions, which further explains the markedly smaller MBC-induced splitting of the $E_{u}$ mode.
+We note that although the $E_{u}$ modes involves vibrations from both Co and S atoms, the atom-resolved MBC shows that S atom vibrations dominate its contribution (details are in the Supplementary Material Fig.
+LABEL:fig:Eu_seperate).
+Combined with Figs.
+2(d2) and  2(e2), it can be observed that while the parity of the $E_{g/u}$ modes remains unchanged at both $\Gamma$ and T, the relative phase between the two S atoms shifts by $\pi$.
+For instance, the $E_{g}$ mode exhibits a relative phase of $\pi$ at $\Gamma$ but 0 at T, whereas the $E_{u}$ mode shows a phase of 0 at $\Gamma$ and $\pi$ at T.
+The distribution of the maximum phonon splitting for the $E_{g/u}$ modes switches along the $\Gamma$–T path, a reversal that correlates with the relative phase between two S atoms of the modes.
+Previous studies have attributed $\Gamma$-point phonon splitting under external magnetic fields to the manifestation of an effective phonon magnetic moment ($M_{ph}$[Niu2021_phononMag, Xue2025_Extrinsic, CeCl3_2022, Bonini2023_CrI3_TRS].
 As detailed in Ref.
-[3], $M_{z}$ comprises two contributions: one from the Born effective charges and another from the Berry curvature of electrons in both real space (MBC) and momentum space ($\Omega_{e}$), i.e.,
-$M_{z}=\frac{e}{2M_{\kappa}}l_{z,\nu\bm{q}}\int d\bm{k}\left\{\Omega_{k_{x}u_{y}}-\Omega_{k_{y}u_{x}}+\Omega_{e,k_{x}k_{y}}G_{u_{x}u_{y}}\right\}.$ (4)
+[Niu2021_phononMag], $M_{ph,z}$ comprises two contributions: one from the Born effective charges and the other from the Berry curvature of electrons in both nuclear coordinates (MBC) and momentum space ($\Omega_{e}$), i.e.,
+$M_{ph,z}=\frac{e}{2M_{\kappa}}l_{z,\nu\bm{q}}\int d\bm{k}\left\{\Omega_{k_{x}u_{y}}-\Omega_{k_{y}u_{x}}+\Omega_{e,k_{x}k_{y}}G_{u_{x}u_{y}}\right\}.$ (4)
 The first two terms collectively integrate to the Born effective charge, whereas the third term is equivalent to the product of the conventional electronic and molecular Berry curvatures.
 In css, the metallic nature and the preservation of $\mathcal{P}$ result in a vanishing Born effective charge.
-Given that $G$ is independent of $\bm{k}$ and the integral $\int d\bm{k}\Omega_{e,k_{x}k_{y}}$ is proportional to the anomalous Hall conductivity $\sigma^{\text{A}}_{xy}$, the expression for the phonon magnetic moment simplifies to $M_{z}\propto l_{z,\nu\bm{q}}\cdot G\cdot\sigma^{\text{A}}_{xy}$.
-Consequently, the large intrinsic Hall conductivity of css (approximately 505 $\Omega^{-1}$ cm-1 [25]) indicates a strong propensity for intrinsic phonon magnetism in this material.
+Given that $G$ is independent of $\bm{k}$ and the integral $\int d\bm{k}\Omega_{e,k_{x}k_{y}}$ is proportional to the anomalous Hall conductivity $\sigma^{\text{A}}_{xy}$, the expression for the phonon magnetic moment simplifies to $M_{ph}\propto l_{z,\nu\bm{q}}\cdot G\cdot\sigma^{\text{A}}_{xy}$.
+Consequently, the large intrinsic Hall conductivity of css (approximately 505 $\Omega^{-1}$ cm-1 [Wang2018_NC]) indicates a strong propensity for intrinsic phonon magnetism in this material.
 As calculated and shown in Fig.
-5, the pronounced difference in the MBC between the $E_{g}$ and $E_{u}$ modes indicates a larger magnetic moment for the $E_{g}$ phonon.
-In summary, our ab initio-MBC methodology captures the electronic order-induced symmetry breaking in the lattice dynamics of css, generalizing the manifestation of phonon magnetism and phonon energy shift effects in magnetic materials.
-Notably, the energy scale of circularly polarized phonons coincides with that of chiral Weyl fermions, suggesting rich opportunities for chiral boson–fermion coupling and enabling control over topological states through phononic and magnetic means [47].
-This efficient and accurate framework for magnetic lattice dynamics is readily extendable to full Brillouin-zone calculations and enables the study of emergent phenomena such as the phonon Hall effect [13, 14, 15, 16, 17, 15, 18, 19, 20, 4], Hall viscosity [48], and so on.
-The algorithm is broadly applicable to intrinsic magnetic systems, facilitating the discovery of materials with large phonon magnetism.
-Our findings establish a set of design principles for materials with large phonon magnetic moments and substantial splitting, i.e., strong spin-orbit coupling, significant electron-phonon coupling, and phonon modes carrying nonzero angular momentum.
-We also propose that materials integrating magnetic order with flat bands, such as css, represent a particularly fertile platform for the discovery of giant phonon splitting and pronounced phonon magnetism.
-Future studies may focus on extrinsic contributions to the phonon magnetic moment, such as relaxation processes [4], non-reciprocal transport related to circularly polarized phonons [49, 21, 50, 51], and ab initio algorithm for external field–induced phonon magnetism, etc.
-I Acknowledgement
+3, the pronounced difference in the MBC between two modes indicates a larger magnetic moment for the $E_{g}$ phonon.
+III Conclusion and Discussion
+In summary, we establish an ab initio MBC framework that captures electronic-order–induced symmetry breaking and phonon magnetism in magnetic materials, exemplified by Co3Sn2S2.
+Remarkably, the energy scale of circularly polarized phonons coincides with that of chiral Weyl fermions, unveiling opportunities for chiral boson–fermion coupling and enabling control of topological states through phononic and magnetic interactions [hernandez2023observation].
+This framework offers an efficient and accurate route to modeling magnetic lattice dynamics, readily extendable to full Brillouin-zone calculations and the exploration of emergent collective phenomena.
+From these results, we derive design principles for engineering large phonon magnetic moments and sizable phonon splitting, which require strong electron–phonon coupling, pronounced SOC, and phonon modes with finite angular momentum.
+Magnetic materials with flat electronic bands, such as Co3Sn2S2, thus provide an ideal platform for realizing giant phonon splitting and pronounced phonon magnetism.
+Beyond Co3Sn2S2, our algorithm also predicts ferromagnetic materials such as FeCo and Co2MnSi, both with $O_{h}$ symmetry, to exhibit FM-order–induced phonon splitting and phonon magnetism in their lattice dynamics.
+The MBC-induced maximum splitting of the $E$ mode along the $C_{4}$-invariant path is calculated to be 1.86 cm-1 for FeCo and 0.16 cm-1 for Co2MnSi.
+Further details are provided in the Supplementary Materials.
+We note that larger splittings may be observed experimentally due to additional effects such as Fano resonances or electronic correlations.
+IV Methods
+The electronic structure and the density functional perturbation theory (DFPT) calculation is performed by the Quantum Espresso package [QE1, QE2, QE3].
+The norm-conserving pseudo-potential of the Perdew-Burke-Ernzerhof (PBE) exchange–correlation functional [perdew_generalized_1996, dojo, ONCV_Hamann] was used.
+The kinetic energy cutoff is set to 90 Ry.
+The Brillouin zone was sampled using a $8\times 8\times 8$ Monkhorst–Pack mesh.
+The MBC is calculated under the first-order-perturbation framework and calculated by the Wannier-based electron-phonon coupling vortex with modified EPW code [mostofi2008wannier90, mostofi2014updated, EPW1, EPW2, EPW3].
+The MBC is Wannier-interpolated on a $80\times 80\times 80$ fine $k$-mesh.
+The core code will be published along with the manuscript.
+Acknowledgments
 We acknowledge the helpful discussion with Haijun Zhang, Fuyi Wang, Xi Dai, Yang Gao, Yafei Ren, Shang Ren, Run Yang and Martin Dressel.
-We acknowledge the support from National Key R&D Project (Grant Nos.
-2023YFA1407400 and 2024YFA1409200), and the National Natural Science Foundation of China Grant No.
-12374165.
+Foundings:
+We acknowledge the support from National Key R&D Project (Grant Nos. 2023YFA1407400 and 2024YFA1409200), and the National Natural Science Foundation of China Grant No. 12374165.
+Author contributions:
+T.Z. conceived of the presented idea. T.Z. and S.Z. developed the theory and performed the first-principles calculations. All authors discussed the results and contributed on the manuscript.
+Competing interests:
+The authors declare that they have no competing interests.
+Data availability:
+All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. The code is available from the corresponding author upon reasonable request.